'two teachers have classes of similar sizes. After final exams, the mean of the grades in each classes in 73%. However, one class has a standard deviation of 4% while the other's is 8%. In which class would a mark of 90% be more meaningful? Explain your reasoning.'

Answer by stanbon(75887)

'two teachers have classes of similar sizes. After final exams, the mean of the grades in each classes in 73%. However, one class has a standard deviation of 4% while the other's is 8%. In which class would a mark of 90% be more meaningful? Explain your reasoning.'
----
Compare the z-scores.
(90-73)/4 = 17/4 = 4.25 standard deviations above the mean
---------
(90-73)/8 = 17/8 = 2.625 standard deviations above the mean
-------------
Cheers,
Stan H.
--------------